# Non-Convex Self-Concordant Functions: Practical Algorithms and Complexity Analysis

### v1

@@ 5 Experiments
Table 1: Test accuracy of different neural network optimizers. We replicate the KFAC experiments from [Zhang-2019-Three] using 10 random seeds. Values following $\pm$” denote standard deviations across seeds.
Dataset Model SGD Adam [Kingma-2015-Adam] KFAC KFAC-WSC
@@
The network is trained by minimizing $\mathcal{L}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}}[L(f_{\theta}(x),y)]$, where $L$ is a loss function, and the FIM is $F(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\nabla_{\theta}\log p_{\theta}(y\mid x)(\nabla_{\theta}\log p_{\theta}(y\mid x))^{\top}]$ where $p_{\theta}(y\mid x)$ denotes the model’s predictive distribution. For the weight matrix $W_{l}$, the gradient of $\mathcal{L}$ with respect to $W_{l}$ can be written in terms of the input activation and the back-propagated signal. In particular, $\nabla_{W_{l}}L=g_{l}a_{l}^{\top}$, where $g_{l}=\nabla_{s_{l}}L\in\mathbb{R}^{d_{\textnormal{out}}}$ is the gradient of $L$ with respect to the pre-activations $s$. The Fisher block corresponding to $W_{l}$ is $F_{l}=\mathbb{E}[\operatorname{vec}(\nabla_{W_{l}}L)\operatorname{vec}(\nabla_{W_{l}}L)^{\top}]=\mathbb{E}[(a_{l}\otimes g_{l})(a_{l}\otimes g_{l})^{\top}]$. The direct computation of $F_{l}$ is intractable, as it scales with the square of the number of parameters in $W$. The key idea behind KFAC is to treat $a_{l}$ and $g_{l}$ as independent, yielding $F_{l}\approx\mathbb{E}[a_{l}a_{l}^{\top}]\otimes\mathbb{E}[g_{l}g_{l}^{\top}]$. For any $X\in\mathbb{R}^{d_{\textnormal{out}}\times d_{\textnormal{in}}}$, we have $F^{-1}_{l}\mathrm{vec}(X)\approx(\mathbb{E}[a_{l}a_{l}^{\top}])^{-1}X(\mathbb{E}[g_{l}g_{l}^{\top}])^{-1}=A_{l}^{-1}XS_{l}^{-1}$, where $A_{l}=\mathbb{E}[a_{l}a_{l}^{\top}]$ and $S_{l}=\mathbb{E}[g_{l}g_{l}^{\top}]$. Thus, if the partial gradient of $\mathcal{L}$ with respect to $W_{l}$ is denoted as $\nabla_{W_{l}}\mathcal{L}(W)$ (as is common in PyTorch), the natural gradient is given by $A_{l}^{-1}\nabla_{W_{l}}\mathcal{L}(W)S_{l}^{-1}$.
When implementing KFAC, it is common to add regularization to $A_{l}$ and $S_{l}$ to improve both generalization and numerical stability [Martens-2015-Optimizing]. The natural gradient with Tikhonov regularization [Martens-2015-Optimizing] is $D_{l}=(A_{l}+\gamma I)^{-1}\nabla_{W_{l}}\mathcal{L}(W)(S_{l}+\gamma I)^{-1}$. We can interpret $D_{l}$ as an approximate Newton direction $(\nabla^{2}(f+F)(x_{j}))^{-1}\nabla f(x_{j})$ in the sense that $\nabla^{2}F(x_{j})$ is replaced by $\gamma I$. We interpret $\gamma$ as the weak self-concordance parameter $\ell$ and $\lambda_{l}=\sqrt{\mathrm{vec}(D_{l})^{\top}\mathrm{vec}(\nabla_{W_{l}}\mathcal{L}(W))}$ as an approximation of $\nu_{f,F}(x_{k})$.
We follow the techniques of [Zhang-2019-Three] and evaluate our algorithms by training VGG16 [Simonyan-2015-Very] and ResNet32 [He-2016-Deep, Zagoruyko-2016-Wide] on CIFAR-10 and CIFAR-100 [Krizhevsky-2009-Learning]. We conduct experiments on a single NVIDIA V100 SXM2 GPU from San Diego Supercomputer Center [Strande-2021-Expanse]. The results are summarized in Table 1. We performed a grid search to identify the best hyperparameters that maximize test accuracy during training. Both KFAC and KFAC-WSC were trained for 100 epochs. The initial learning rate was selected from the set $\eta_{0}\in\{10^{-3},3\times 10^{-2},10^{-2}\}$ and the learning rate decay by $0.1$ was applied at the 40th and 80th epochs. The weight decay factor from $\omega\in\{10^{-2},3\times 10^{-1},10^{-1}\}$, and the damping factor from $\gamma\in\{3\times 10^{-3},10^{-2},3\times 10^{-2}\}$. Our method improved training efficiency on VGG16 for CIFAR-10 and CIFAR-100, while being outperformed by KFAC on ResNet32.
@@ 6 Conclusions
We generalized the notion of self-concordance to non-convex settings by introducing weakly self-concordant functions and $F$-based self-concordant functions. Based on these two classes, we proposed regularized Newton and adaptive regularization methods and proved their global and local convergence guarantees. We also conducted numerical experiments to evaluate their effectiveness. Indeed, our adaptive regularization method outperformed adaptive cubic regularization, L-BFGS, and trust-region methods on non-negative matrix factorization problems in terms of both MSE and KL divergence losses. Our KFAC-based regularized Newton method improved training efficiency on VGG16 for CIFAR-10 and CIFAR-100, while being outperformed by KFAC on ResNet32. In summary, our results demonstrate the importance of extending self-concordance for large-scale non-convex optimization.
@@ Acknowledgments
The second author is partially supported by a start-up fund at the University of Hong Kong and by the Hong Kong Research Grants Council under ECS project 27301425. The third author is partially supported by a start-up fund at Columbia University. We also thank the support of GPU hours from NSF ACCESS [Boerner-2023-Access].
@@ A.1 Convergence to a Proximal Stationary Point
For any $\eta>0$, we define the Moreau envelope $f_{\eta}:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ [Moreau-1965-Proximite] of $f$ by $f_{\eta}(y)=\inf_{x\in\mathbb{R}^{n}}\{f(x)+\tfrac{1}{2\eta}\|y-x\|^{2}\}$. The gradient norms of the Moreau envelope, $\|\nabla f_{\eta}(x)\|$, have been used as a measure of stationarity for weakly convex functions [Davis-2019-Stochastic]. For weakly self-concordant functions, we show next that RNM (see Eq. (8)) and Algorithm 1 can find a point with $\|\nabla f_{\eta}(x)\|\leq\epsilon$ (for any small enough $\eta$) in $O(\epsilon^{-2})$ iterations.
@@
Let $\epsilon^{\prime}=\frac{\sqrt{\mu-\ell}\epsilon}{\mu+\kappa\sqrt{\mu-\ell}\epsilon}$ and $\varphi(y)=f(y)+\frac{1}{2}\mu\|y-x\|^{2}$. Thus, we assume $\nu_{f,\frac{1}{2}\mu\|\cdot\|^{2}}(x)=\lambda_{\varphi}(x)\leq\epsilon^{\prime}$. It follows from [Davis-2019-Stochastic, Lemma 2.2] that $\tfrac{1}{\mu}{\|\nabla f_{1/\mu}(x)\|}=\|x-\arg\min\varphi\|$. Moreover, $\nabla^{2}\varphi\geq(\mu-\ell)I$ implies that $\sqrt{\mu-\ell}\|x-\arg\min\varphi\|\leq\|x-\arg\min\varphi\|_{\nabla^{2}\varphi(x)}$. We define $\omega_{\star}^{\prime}(t)=\frac{t}{1-t}$ for $t\in[0,1)$ and apply [Nesterov-2018-Lectures, Eq. (5.2.4)] with $f=\varphi$. In the notation of [Nesterov-2018-Lectures], we note that $x_{f}^{\star}=\arg\min\varphi$, $M_{f}=\kappa$, $\|\cdot\|_{x}=\|\cdot\|_{\nabla^{2}\varphi(x)}$. Thus, we have $\|x-\arg\min_{y}\varphi(y)\|_{\nabla^{2}\varphi(x)}\leq\kappa^{-1}\omega_{\star}^{\prime}(\kappa\lambda_{\varphi}(x))$. Therefore, we obtain $\frac{\sqrt{\mu-\ell}}{\mu}\|\nabla f_{1/\mu}(x)\|\leq\|x-\arg\min\varphi\|_{\nabla^{2}\varphi(x)}\leq\kappa^{-1}\omega_{\star}^{\prime}(\kappa\lambda_{\varphi}(x))\leq\kappa^{-1}\omega_{\star}^{\prime}(\kappa\epsilon^{\prime})=\tfrac{\sqrt{\mu-\ell}}{\mu}\epsilon$. ∎
@@ A.2 Inexact Proximal Point Method (IPPM)
We develop an algorithm for minimizing a $(\kappa,\ell)$-weakly self-concordant function $f$ using only gradient and Hessian–vector product (HVP) oracles. Our approach is based on the proximal point method [rockafellar1976monotone]. Given the current iterate $z_{j}\in\textbf{dom}(f)$, we consider the regularized objective $f_{j}:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ defined as $f_{j}(y)=f(y)+\frac{1}{2}\mu\|y-z_{j}\|^{2}$. The update of the proximal point method is given by $z_{j+1}=\arg\min f_{j}$. When $\mu>\ell$, each function $f_{j}$ is self-concordant.
To solve these subproblems inexactly, we introduce Algorithm 3. The algorithm combines the Newton–CG (conjugate gradient) method [Wright-2006-Numerical, Algorithm 7.1] with the Lanczos method [Kuczynski-1992-Estimating]. By exploiting self-concordance, it avoids line searches, which is its main advantage over traditional Newton–CG. For a self-concordant function $f$, the algorithm outputs, with probability $1-\delta$, $y\in\textbf{dom}(f)$ such that either $\lambda_{f}(y)\leq\epsilon_{1}$ and $\lambda_{f}(x_{0})\geq\sqrt{\frac{1-\alpha_{\star}}{1+\alpha_{\star}}}\epsilon_{0}$, or $y=x_{0}$ and $\lambda_{f}(x_{0})\leq\epsilon_{0}$. A more precise convergence theorem, together with its proof, is deferred to Theorem A.10 in Appendix A.3. This algorithm may be of independent interest in convex optimization.
@@
6:  $\rho_{k}\leftarrow h_{k}^{\top}g_{k}$, $\delta_{k}\leftarrow{\nabla^{2}f(x_{k})[h_{k},h_{k}]}$.
7:  return $x_{k}$ if $\rho_{k}\leq(1-\alpha_{\star})\epsilon_{1}^{2}$ or ($k=0$ and $\rho_{0}\leq(1-\alpha_{\star})\epsilon_{0}^{2}$).
8:  if $\rho_{k}>\frac{R_{1}^{2}}{\kappa^{2}}$ then
9:   $t_{k}\leftarrow\frac{\rho_{k}}{\delta_{k}+\kappa\rho_{k}\sqrt{\delta_{k}}}$, $B_{k}\leftarrow(1+\frac{\sqrt{1+\alpha_{\star}}}{1-\alpha_{\star}}\kappa\sqrt{\rho_{k}})^{2}$, $\beta_{k+1}\leftarrow B_{k}\beta_{k}$.
10:  else
11:   if flag then
12:    $\beta^{\star}\leftarrow{\text{sqrt-cond}(\nabla^{2}f(x_{k}),0.5\delta,\beta_{k})}$, $\beta_{\mathrm{local}}\leftarrow C_{3}^{4}\beta^{\star}$, $\texttt{flag}\leftarrow\texttt{False}$.
13:   end if
14:   $t_{k}\leftarrow\frac{\rho_{k}}{\delta_{K}+2\kappa\rho_{k}\sqrt{\delta_{k}}}$, $\beta_{k+1}\leftarrow\beta_{\mathrm{local}}$.
15:  end if
16:  $x_{k+1}\leftarrow x_{k}-t_{k}h_{k}$.
17:end for
18:function sqrt-cond($H,\delta,\beta$):
19:    Use the Lanczos algorithm to obtain $\lambda_{1},\lambda_{2}$ such that $|\lambda_{1}-\lambda_{\max}(H)|\leq\frac{1}{3}\lambda_{\max}(H)$ and $|\lambda_{2}-\lambda_{\min}(H)|\leq\frac{1}{3}\lambda_{\min}(H)$. return $\sqrt{2\lambda_{1}/\lambda_{2}}$.
20:function CG-Inverse($H,g,\beta,\alpha$):
21:    Perform $\min\left\{n,\left\lfloor\log_{\tfrac{\beta-1}{\beta+1}}{0.5\alpha}\right\rfloor+1\right\}$ CG iterations to compute $H^{-1}g$ start from 0.
@@
6:  $z_{j+1}\leftarrow$ The output of Algorithm 3 with objective function $f_{j}$, input $\kappa\leftarrow\kappa$, $\epsilon_{0}\leftarrow\epsilon$, $\epsilon_{1}\leftarrow\epsilon^{\prime}$, $\beta\leftarrow\beta^{\prime}$, $\delta\leftarrow\delta^{\prime}$, and initialization $x_{0}\leftarrow z_{j}$.
7:  return $z_{j}$ if $z_{j+1}=z_{j}$.
8:end for
@@
In IPPM, there exists $B>0$ such that for all $j\geq 0$, we have $\mathrm{cond}(\nabla^{2}f(z_{0})+\mu I)\leq B^{2}$ and $\mathrm{cond}(\nabla^{2}f_{j}(\arg\min f_{j})+\mu I)\leq B^{2}$.

### v2

@@ 4 Experiments
The network is trained by minimizing $\mathcal{L}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}}[L(f_{\theta}(x),y)]$, where $L$ is a loss function, and the FIM is $P(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\nabla_{\theta}\log p_{\theta}(y\mid x)(\nabla_{\theta}\log p_{\theta}(y\mid x))^{\top}]$ where $p_{\theta}(y\mid x)$ denotes the model’s predictive distribution. For the weight matrix $W_{l}$, the gradient of $\mathcal{L}$ with respect to $W_{l}$ can be written in terms of the input activation and the back-propagated signal. In particular, $\nabla_{W_{l}}L=g_{l}a_{l}^{\top}$, where $g_{l}=\nabla_{s_{l}}L\in\mathbb{R}^{d_{\textnormal{out}}}$ is the gradient of $L$ with respect to the pre-activations $s$. The Fisher block corresponding to $W_{l}$, equivalently, the diagonal block of the Gauss–Newton matrix, is given by $P_{l}=\mathbb{E}[\operatorname{vec}(\nabla_{W_{l}}L)\operatorname{vec}(\nabla_{W_{l}}L)^{\top}]=\mathbb{E}[(a_{l}\otimes g_{l})(a_{l}\otimes g_{l})^{\top}]$. The direct computation of $P_{l}$ is intractable, as it scales with the square of the number of parameters in $W$. The key idea behind KFAC is to treat $a_{l}$ and $g_{l}$ as independent, yielding $P_{l}\approx\mathbb{E}[a_{l}a_{l}^{\top}]\otimes\mathbb{E}[g_{l}g_{l}^{\top}]$. For any $X\in\mathbb{R}^{d_{\textnormal{out}}\times d_{\textnormal{in}}}$, we have $P^{-1}_{l}\mathrm{vec}(X)\approx(\mathbb{E}[a_{l}a_{l}^{\top}])^{-1}X(\mathbb{E}[g_{l}g_{l}^{\top}])^{-1}=A_{l}^{-1}XS_{l}^{-1}$, where $A_{l}=\mathbb{E}[a_{l}a_{l}^{\top}]$ and $S_{l}=\mathbb{E}[g_{l}g_{l}^{\top}]$. Consequently, denoting the partial gradient of $\mathcal{L}$ with respect to $W_{l}$ as $\nabla_{W_{l}}\mathcal{L}(W)\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$, the natural gradient (i.e., the Gauss–Newton update) is approximated by $A_{l}^{-1}\nabla_{W_{l}}\mathcal{L}(W)\,S_{l}^{-1}$.
In practice, it is common to add regularization to $A_{l}$ and $S_{l}$ to improve both generalization and numerical stability [47]. The Gauss-Newton update with Tikhonov regularization [47] is $D_{l}=-(A_{l}+\gamma I)^{-1}\nabla_{W_{l}}\mathcal{L}(W)(S_{l}+\gamma I)^{-1}$. We interpret $D_{l}$ as an instance of the regularized Newton direction $d_{j}=-(\nabla^{2}(f+F)(x_{j}))^{-1}\nabla f(x_{j})$ from Algorithm 1, where the Hessian regularization term $\nabla^{2}F(x_{j})$ is replaced by the scalar shift $\gamma I$, and the damping parameter $\gamma$ plays the role of the weak self-concordance parameter $\ell$. Accordingly, we identify $\lambda_{l}=\sqrt{-\operatorname{vec}(D_{l})^{\top}\operatorname{vec}(\nabla_{W_{l}}\mathcal{L}(W))}$ as a layer-wise approximation of the quantity $\lambda_{j}=\sqrt{-\nabla f(x_{j})^{\top}d_{j}}$ computed in Line 4 of Algorithm 1—namely, the square root of the inner product between the gradient and the negative update direction, which serves as the regularized Newton decrement. Our KFAC-WSC implementation (Algorithm 3) therefore shares the same underlying structure as Algorithm 1 and the classical damped Newton method for convex self-concordant functions [58]: the main difference over regular Gauss-Newton’s method is that the step size applied to the (regularized) Newton update is taken proportional to $\frac{1}{1+\kappa\lambda_{l}}$, where $\lambda_{l}$ is the square root of the regularized Newton decrement, as prescribed by Line 5 of Algorithm 1. Consequently, the step sizes in KFAC-WSC are different for different layers.
We follow the techniques of [82] and evaluate our algorithms by training VGG16 [69] and ResNet32 [35, 80] on CIFAR-10 and CIFAR-100 [37]. The test accuracies of SGD and Adam [36] are copied from [82], where a grid search was conducted to find the best hyperparameters. For KFAC [47] and KFAC-WSC (our method), we conducted our own experiments on a single NVIDIA V100 SXM2 GPU from San Diego Supercomputer Center [70]. The results are summarized in Table 1. Both KFAC and KFAC-WSC were trained for 100 epochs, and the learning rate was decayed by a factor of $0.1$ at the 40th and 80th epochs. We performed grid searches to identify the hyperparameters that maximize test accuracy during training. For KFAC, we adopted the best hyperparameters from [82], on top of which we tuned the learning rate over $\{3\times 10^{-4},1\times 10^{-3},3\times 10^{-3}\}$, selecting the best learning rate based on validation accuracy. Our results match [82] closely. For KFAC-WSC, the initial learning rate was selected from $\eta_{0}\in\{10^{-3},3\times 10^{-3},10^{-2}\}$. The weight decay factor was chosen from $\omega\in\{10^{-2},3\times 10^{-2},10^{-1},3\times 10^{-1},1.0\}$, and the damping factor from $\gamma\in\{3\times 10^{-3},10^{-2},3\times 10^{-2}\}$. This setup matches the $3\times 5\times 3$ grid used in [78], a PyTorch implementation of [82]. In practice, using the same hyperparameters as KFAC for KFAC-WSC leads to degraded performance. This is evident since we are using smaller learning rates for each layer. Our method improved training efficiency on VGG16 for CIFAR-10 and CIFAR-100, while being outperformed by KFAC on ResNet32.
A convergence analysis of KFAC [47] and, its extension, KFAC-WSC is beyond the scope of this paper. Rather, the goal of this experiment is to empirically assess the effectiveness of incorporating the self-concordant step size rule from Algorithm 1 into a practical second-order optimizer.
@@
Table 1: Test accuracy of different neural network optimizers. We replicate the KFAC experiments from [82] using 10 random seeds and copied the results of SGD and Adam ([36, 43]) from [82]. Values following $\pm$” denote standard deviations across seeds. $(\eta_{0}^{*},\omega^{*},\gamma^{*})$ are the best used initial learning rate, weight decay, and damping factor respectively selected from grid searches.
SGD Adam KFAC KFAC-WSC
Dataset Model Acc. Acc. Acc. $(\eta_{0}^{*},\omega^{*},\gamma^{*})$ Acc. $(\eta_{0}^{*},\omega^{*},\gamma^{*})$
@@ 5 Conclusions
We generalized the notion of self-concordance to non-convex settings by introducing weakly self-concordant functions and $F$-based self-concordant functions. Based on these two classes, we proposed regularized Newton and adaptive regularization methods and proved their global and local convergence guarantees. We also conducted a few preliminary numerical experiments to evaluate their effectiveness. Indeed, our adaptive regularization method outperformed adaptive cubic regularization, L-BFGS, and trust-region methods on non-negative matrix factorization problems in terms of both MSE and KL divergence losses. Our KFAC-based regularized Newton method improved training efficiency on VGG16 for CIFAR-10 and CIFAR-100, while being outperformed by KFAC on ResNet32. In summary, we believe that the analytic and experimental results that we presented illustrate the potential of extending the powerful properties of self-concordance to non-convex optimization.
@@ Acknowledgments
The second author is partially supported by a start-up fund at the University of Hong Kong and by the Hong Kong Research Grants Council under ECS project 27301425. The third author is partially supported by a start-up fund at Columbia University. We also thank the support of GPU hours from NSF ACCESS [7].
@@ A.1 Convergence to a Proximal Stationary Point
For any $\eta>0$, we define the Moreau envelope $f_{\eta}:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ [54] of $f$ by $f_{\eta}(y)=\inf_{x\in\mathbb{R}^{n}}\{f(x)+\tfrac{1}{2\eta}\|y-x\|^{2}\}$. The gradient norms of the Moreau envelope, $\|\nabla f_{\eta}(x)\|$, have been used as a measure of stationarity for weakly convex functions [21]. For weakly self-concordant functions, we show next that RNM (see Algorithm 1) and Algorithm 2 can find a point with $\|\nabla f_{\eta}(x)\|\leq\epsilon$ (for any small enough $\eta$) in $O(\epsilon^{-2})$ iterations.
@@
Proof Let $\epsilon^{\prime}=\frac{\sqrt{\mu-\ell}\epsilon}{\mu+\kappa\sqrt{\mu-\ell}\epsilon}$ and $\varphi(y)=f(y)+\frac{1}{2}\mu\|y-x\|^{2}$. Thus, we assume $\nu_{f,\frac{1}{2}\mu\|\cdot\|^{2}}(x)=\lambda_{\varphi}(x)\leq\epsilon^{\prime}$. It follows from [21, Lemma 2.2] that $\tfrac{1}{\mu}{\|\nabla f_{1/\mu}(x)\|}=\|x-\arg\min\varphi\|$. Moreover, $\nabla^{2}\varphi\geq(\mu-\ell)I$ implies that $\sqrt{\mu-\ell}\|x-\arg\min\varphi\|\leq\|x-\arg\min\varphi\|_{\nabla^{2}\varphi(x)}$. We define $\omega_{\star}^{\prime}(t)=\frac{t}{1-t}$ for $t\in[0,1)$ and apply [58, Eq. (5.2.4)] with $f=\varphi$. In the notation of [58], we note that $x_{f}^{\star}=\arg\min\varphi$, $M_{f}=\kappa$, $\|\cdot\|_{x}=\|\cdot\|_{\nabla^{2}\varphi(x)}$. Thus, we have $\|x-\arg\min_{y}\varphi(y)\|_{\nabla^{2}\varphi(x)}\leq\kappa^{-1}\omega_{\star}^{\prime}(\kappa\lambda_{\varphi}(x))$. Therefore, we obtain $\frac{\sqrt{\mu-\ell}}{\mu}\|\nabla f_{1/\mu}(x)\|\leq\|x-\arg\min\varphi\|_{\nabla^{2}\varphi(x)}\leq\kappa^{-1}\omega_{\star}^{\prime}(\kappa\lambda_{\varphi}(x))\leq\kappa^{-1}\omega_{\star}^{\prime}(\kappa\epsilon^{\prime})=\tfrac{\sqrt{\mu-\ell}}{\mu}\epsilon$. $\square$
@@ A.2 Inexact Proximal Point Method (IPPM)
We develop an algorithm for minimizing a $(\kappa,\ell)$-weakly self-concordant function $f$ using only gradient and Hessian–vector product (HVP) oracles. Our approach is based on the proximal point method [64]. Given the current iterate $z_{j}\in\textbf{dom}(f)$, we consider the regularized objective $f_{j}:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ defined as $f_{j}(y)=f(y)+\frac{1}{2}\mu\|y-z_{j}\|^{2}$. The update of the proximal point method is given by $z_{j+1}=\arg\min f_{j}$. When $\mu>\ell$, each function $f_{j}$ is self-concordant.
To solve these subproblems inexactly, we introduce Algorithm 4. The algorithm combines the Newton–CG (conjugate gradient) method [60, Algorithm 7.1] with the Lanczos method [38]. By exploiting self-concordance, it avoids line searches, which is its main advantage over traditional Newton–CG. For a self-concordant function $f$, the algorithm outputs, with probability $1-\delta$, $y\in\textbf{dom}(f)$ such that either $\lambda_{f}(y)\leq\epsilon_{1}$ and $\lambda_{f}(x_{0})\geq\sqrt{\frac{1-\alpha_{\star}}{1+\alpha_{\star}}}\epsilon_{0}$, or $y=x_{0}$ and $\lambda_{f}(x_{0})\leq\epsilon_{0}$. A more precise convergence theorem, together with its proof, is deferred to Theorem A.10 in Appendix A.3. This algorithm may be of independent interest in convex optimization.
@@
6:  $\rho_{k}\leftarrow-h_{k}^{\top}g_{k}$, $\delta_{k}\leftarrow\nabla^{2}f(x_{k})[h_{k},h_{k}]$. return $x_{k}$ if $\rho_{k}\leq(1-\alpha_{\star})\epsilon_{1}^{2}$ or ($k=0$ and $\rho_{0}\leq(1-\alpha_{\star})\epsilon_{0}^{2}$).
7:  if $\rho_{k}>\frac{R_{1}^{2}}{\kappa^{2}}$ then
8:    $t_{k}\leftarrow\frac{\rho_{k}}{\delta_{k}+\kappa\rho_{k}\sqrt{\delta_{k}}}$, $B_{k}\leftarrow(1+\frac{\sqrt{1+\alpha_{\star}}}{1-\alpha_{\star}}\kappa\sqrt{\rho_{k}})^{2}$, $\beta_{k+1}\leftarrow B_{k}\beta_{k}$.
9:  else
10:    if flag then
11:     $\beta^{\star}\leftarrow\text{sqrt-cond}(\nabla^{2}f(x_{k}),0.5\delta,\beta_{k})$, $\beta_{\mathrm{local}}\leftarrow C_{3}^{4}\beta^{\star}$, $\texttt{flag}\leftarrow\texttt{False}$.
12:    end if
13:    $t_{k}\leftarrow\frac{\rho_{k}}{\delta_{k}+2\kappa\rho_{k}\sqrt{\delta_{k}}}$, $\beta_{k+1}\leftarrow\beta_{\mathrm{local}}$.
14:  end if
15:  $x_{k+1}\leftarrow x_{k}+t_{k}h_{k}$.
16:end for
17:function sqrt-cond($H,\delta,\beta$):
18:   Use the Lanczos algorithm to obtain $\lambda_{1},\lambda_{2}$ such that $|\lambda_{1}-\lambda_{\max}(H)|\leq\frac{1}{3}\lambda_{\max}(H)$ and $|\lambda_{2}-\lambda_{\min}(H)|\leq\frac{1}{3}\lambda_{\min}(H)$. return $\sqrt{2\lambda_{1}/\lambda_{2}}$.
19:function CG-Inverse($H,g,\beta,\alpha$):
20:   Perform $\min\left\{n,\left\lfloor\log_{\tfrac{\beta-1}{\beta+1}}{0.5\alpha}\right\rfloor+1\right\}$ CG iterations to compute $H^{-1}g$ start from 0.
@@
6:  $z_{j+1}\leftarrow$ The output of Algorithm 4 with objective function $f_{j}$, input $\kappa\leftarrow\kappa$, $\epsilon_{0}\leftarrow\epsilon$, $\epsilon_{1}\leftarrow\epsilon^{\prime}$, $\beta\leftarrow\beta^{\prime}$, $\delta\leftarrow\delta^{\prime}$, and initialization $x_{0}\leftarrow z_{j}$. return $z_{j}$ if $z_{j+1}=z_{j}$.
7:end for
@@
In IPPM, there exists $B>0$ such that $\mathrm{cond}(\nabla^{2}f(z_{0})+\mu I)\leq B^{2}$ and we have $\mathrm{cond}(\nabla^{2}f_{j}(\arg\min f_{j}))\leq B^{2}$ for all $j\geq 0$.